### v1

@@ 3.3 Noisy Non-uniform Generatability
Similar to the characterization of non-uniform generatability in the noise-free setting, we can use the noisy uniform generatability to provide sufficiency and necessary conditions for noisy non-uniform generatability.
Lemma 3.5 (Sufficiency for Noisy Non-uniform Generatability).
@@
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\cdots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{i}$ and $\operatorname{NC}_{i}(\mathcal{H}_{i})<\infty$ for all $i\in\mathbb{N}$, then $\mathcal{H}$ is noisily non-uniformly generatable.
@@
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any class satisfying the $\operatorname{UUS}$ property. Suppose there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\cdots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{i}$ and $\operatorname{NC}_{i}(\mathcal{H}_{i})<\infty$ for all $i\in\mathbb{N}$. For a sequence $x_{1},x_{2},\dots$, let $d_{t}:=|\{x_{1},\dots,x_{t}\}|$ denote the number of distinct examples amongst the first $t$ examples. Consider the following generator $\mathcal{G}.$ At time $t\in\mathbb{N}$, $\mathcal{G}$ computes
$j_{t}=\max\{i\in\mathbb{N}:\operatorname{NC}_{i}(\mathcal{H}_{i})<d_{t}\}\cup\{0\}.$
If $j_{t}=0,$ $\mathcal{G}$ plays any $x\in\mathcal{X}.$ Otherwise, it plays from $\langle x_{1},\dots,x_{t}\rangle_{\mathcal{H}_{j_{t}},j_{t}}\setminus\{x_{1},x_{2},\dots,x_{t}\}.$ We claim that $\mathcal{G}$ is a noisy non-uniform generator for $\mathcal{H}$. To see why, let $n^{\star}\in\mathbb{N}$ and $h\in\mathcal{H}$ be the noise level and hypothesis picked by the adversary. Let $i^{\star}\in\mathbb{N}$ be the smallest index such that $h\in\mathcal{H}_{i^{\star}}$ and define $j^{\star}=\max\{i^{\star},n^{\star}\}$. We claim that $\mathcal{G}$ perfectly generates after observing $\operatorname{NC}_{j^{\star}}(\mathcal{H}_{j^{\star}})+1$ distinct examples. To that end, let $x_{1},x_{2},\dots$ be any sequence such that
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq n^{\star}.$
Without loss of generality, suppose there exists $t^{\star}\in\mathbb{N}$ such that $d_{t^{\star}}=\operatorname{NC}_{j^{\star}}(\mathcal{H}_{j^{\star}})+1.$ Fix any $s\geq t^{\star}$. By definition of $j_{t}$, it must be the case that $j_{s}\geq j^{\star}\geq n^{\star}.$ Moreover, because $j_{s}\geq i^{\star}$, we also have that $h\in\mathcal{H}_{j_{s}}$. Together, this means that $h\in\mathcal{H}_{j_{s}}(x_{1},\dots,x_{s},j_{s})$ and $\langle x_{1},\dots,x_{s}\rangle_{\mathcal{H}_{j_{s}},j_{s}}\subseteq\operatorname{supp}(h)$. Also, because $d_{s}>\operatorname{NC}_{j_{s}}(\mathcal{H}_{j_{s}})$ it must be the case that $|\langle x_{1},\dots,x_{s}\rangle_{\mathcal{H}_{j_{s}},j_{s}}|=\infty.$ Accordingly, $\langle x_{1},\dots,x_{s}\rangle_{\mathcal{H}_{j_{s}},j_{s}}\setminus\{x_{1},\dots,x_{s}\}\neq\emptyset$ and $\mathcal{G}$ is guaranteed to output an example in $\operatorname{supp}(h)$ on round $s$. Since $s\geq t^{\star}$ was chosen arbitrarily, this is true for all such $s$, completing the proof. ∎
As a corollary of Lemma 3.5 and Corollary 3.3, we can show that all countable classes are noisily non-uniformly generatable, and hence noisily generatable in the limit. In some sense, this result shows that amongst countable classes, noisy generation comes for free!
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If $\mathcal{H}$ is countably infinite, then $\mathcal{H}$ is noisily non-uniformly generatable, and therefore also noisily generatable in the limit.
@@
Since noisy non-uniform generatability implies noisy generatability in the limit, it suffices to show that every countable $\mathcal{H}$ is noisily non-uniformly generatable. Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any countable class satisfying the UUS property. Let $h_{1},h_{2},\dots$ be some fixed enumeration of $\mathcal{H}$ and define $\mathcal{H}_{n}=\{h_{i}:i\leq n\}$ for all $n\in\mathbb{N}$. Then, observe that $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\dots$ and that $\mathcal{H}=\bigcup_{n=1}^{\infty}\mathcal{H}_{n}$. Next, observe that $|\mathcal{H}_{n}|=n<\infty$, which, using Theorem 3.2, implies that $\operatorname{NC}_{n}(\mathcal{H}_{n})<\infty$. Finally, Lemma 3.5 gives that $\mathcal{H}$ is noisily non-uniformly generatable. ∎
Corollary 3.6 with Lemma 3.4 also establishes that noisy uniform generatability is strictly harder than noisy non-uniform generatability. We now move to our necessity condition for noisy non-uniform generatability.
Lemma 3.7 (Necessity for Noisy Non-uniform Generatability).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If $\mathcal{H}$ is noisily non-uniformly generatable, then for every $n\in\mathbb{N}$, there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\cdots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{i}$ and $\operatorname{NC}_{n}(\mathcal{H}_{i})<\infty$ for all $i\in\mathbb{N}.$
Proof.
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any noisily non-uniformly generatable class satisfying the $\operatorname{UUS}$ property. Suppose that $\mathcal{H}$ is not uniformly noisily generatable and $\mathcal{G}$ be a noisy non-uniform generator for $\mathcal{H}$. Fix $n\in\mathbb{N}$. For every $h\in\mathcal{H}$, let $d_{h,n}\in\mathbb{N}$ be the smallest natural number such that for any sequence $x_{1},x_{2},\dots$ with $\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq n,$ if there exists a $t\in\mathbb{N}$ such that |$|\{x_{1},\dots,x_{t}\}|=d_{h,n}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t$. Let $\mathcal{H}_{i}=\{h\in\mathcal{H}:d_{h,n}\leq i\}$ for all $i\in\mathbb{N}$. Note that $\mathcal{H}=\bigcup_{i\in\mathbb{N}}\mathcal{H}_{i}$ because $d_{h,n}<\infty$ for all $h\in\mathcal{H}.$ Moreover, we have that $\mathcal{H}_{i}\subseteq\mathcal{H}_{i+1}$ for all $i\in\mathbb{N}.$ Finally, for every $i\in\mathbb{N}$, observe that $\mathcal{G}$ is a uniform generator for $\mathcal{H}_{i}$ at noise level $n\in\mathbb{N}$, implying that $\operatorname{NC}_{n}(\mathcal{H}_{i})<i<\infty.$
Note the sufficiency and necessary conditions in Lemmas 3.5 and 3.7 respectively are not matching. Indeed, the condition in Lemma 3.5 is stronger than that in Lemma 3.7. We leave as an open question a complete characterization of noisy non-uniform generatability.
@@ 3.4 Noisy Generatability in the Limit
Corollary 3.6 showed that all countable classes are noisily generatable in the limit. Here, we provide alternate sufficiency conditions for noisy generatability in the limit. Our first result shows that (noiseless) non-uniform generatability is sufficient for noisy generatability in the limit. Since all countable classes are (noiseless) non-uniform generatable (Li et al., 2024), this result also shows that all countable classes noisily generatable in the limit.
@@
1
2for $t=1,2,\dots$ do
3
4      Adversary reveals example $x_{t}$
@@
5      Let $d_{t}=|\{x_{1},\dots,x_{t}\}|$ and $r_{t}\leq t$ be the largest time point such that |$|\{x_{r_{t}},\dots,x_{t}\}|=\lfloor\frac{d_{t}}{2}\rfloor$
6      Initialize $\hat{z}^{t}_{1}=\mathcal{Q}(x_{r_{t}:t})$
7      for $i=1,\dots,r_{t}$ do
8
9            if $\hat{z}^{t}_{i}\notin\{x_{1},\dots,x_{t}\}$ then
10                   $\mathcal{G}$ outputs $\hat{z}^{t}_{i}$ and moves to next round.
11             else
12                   Update $\hat{z}^{t}_{i+1}=\mathcal{Q}(x_{r_{t}},\dots,x_{t},\hat{z}^{t}_{1},\dots,\hat{z}^{t}_{i})$
13
14       end for
15      $\mathcal{G}$ outputs $\hat{z}^{t}_{r_{t}+1}$ and moves to next round.
16 end for
Algorithm 1 Generator $\mathcal{G}$
@@ 6      Initialize $\hat{z}^{t}_{1}=\mathcal{Q}(x_{r_{t}:t})$
We will show that $\mathcal{G}$ noisily generates from $\mathcal{H}$ in the limit. Let $h^{\star}\in\mathcal{H}$ and $x_{1},x_{2},\dots$ be the chosen hypothesis and noisy enumeration of $\operatorname{supp}(h^{\star})$ picked by the adversary, respectively. Let $d_{\mathcal{Q}}(h^{\star})\in\mathbbm{N}$ be the number of noiseless distinct elements that $\mathcal{Q}$ needs for $h^{\star}$ to perfectly generate from $\operatorname{supp}(h^{\star})$. Let $t^{\star}\in\mathbb{N}$ be such that for all $t\geq t^{\star}$, we have that $x_{t}\in\operatorname{supp}(h^{\star}).$ Such a $t^{\star}$ must exist because there are most a finite number of noisy examples in the stream. Also, because $x_{1},x_{2},\dots$ is a noisy enumeration, at some point $s^{\star}\in\mathbb{N}$, $r_{s^{\star}}\geq t^{\star}$ and $\lfloor\frac{d_{s^{\star}}}{2}\rfloor\geq d_{\mathcal{Q}}(h^{\star})$. Fix some $s\geq s^{\star}.$ We will prove that $\mathcal{G}$ generates perfectly on round $s$. First, we claim that for all $i\in[r_{s}+1]$, we have that $\hat{z}^{s}_{i}\in\operatorname{supp}(h^{\star})\setminus\{x_{r_{t}},\dots,x_{t},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{i-1}\}$. Our proof will be by induction. For the base case, consider $i=1$. We need to show that $\hat{z}^{s}_{1}\in\operatorname{supp}(h^{\star})\setminus\{x_{r_{t}},\dots,x_{t}\}$. However, this just follows from the fact that $\hat{z}^{s}_{1}=\mathcal{Q}(x_{r_{t}:t})$, |$|\{x_{r_{t}},\dots,x_{t}\}|\geq d_{\mathcal{Q}}(h^{\star})$, and $\{x_{r_{t}},\dots,x_{t}\}\subseteq\operatorname{supp}(h^{\star}).$ Next, for the induction step, suppose that for all $i\leq m$, we have that $\hat{z}^{s}_{i}\in\operatorname{supp}(h)\setminus\{x_{r_{t}},\dots,x_{t},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{i}\}$. Then, by definition, we have that $\hat{z}_{m+1}^{s}=\mathcal{Q}(x_{r_{t}},\dots,x_{t},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{m})$. The proof of the claim is complete after noting that $\{x_{r_{t}},\dots,x_{t},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{m}\}\subseteq\operatorname{supp}(h^{\star})$ and |$|\{x_{r_{t}},\dots,x_{t},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{m}\}|\geq d_{\mathcal{Q}}(h^{\star}).$
@@
Now we will complete the overall proof by showing that the output of $\mathcal{G}$ on round $s$ lies in $\operatorname{supp}(h^{\star})\setminus\{x_{1},\dots,x_{t}\}$. Suppose that $\mathcal{G}$ outputs $\hat{z}_{j}^{s}$ for some $j\in[r_{t}+1]$. If $j<r_{t}+1$, then it must be the case that Line 7 fired and so by construction $\hat{z}_{j}^{s}\in\operatorname{supp}(h^{\star})\setminus\{x_{1},\dots,x_{t}\}$. Thus, suppose that $j=r_{t}+1.$ If $j=r_{t}+1$, then it must be the case that $\{\hat{z}_{1}^{s},\dots,\hat{z}_{r_{t}}^{s}\}=\{x_{1},\dots,x_{r_{t}}\}.$ Accordingly, we have that $\{x_{r_{t}},\dots,x_{t},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{r_{t}}\}=\{x_{1},\dots,x_{t}\}$, which means that $\hat{z}^{s}_{r_{t}+1}\in\operatorname{supp}(h^{\star})\setminus\{x_{r_{t}},\dots,x_{t},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{r_{t}}\}$ and therefore $\hat{z}^{s}_{r_{t}+1}\in\operatorname{supp}(h^{\star})\setminus\{x_{1},\dots,x_{t}\}.$ Thus, in either case, $\mathcal{G}$ perfectly generates on round $s$. Since $s\geq s^{\star}$ is arbitrary, our proof is complete. ∎
@@
In similar spirit to Theorem 3.10 from Li et al. (2024), we provide a another sufficiency condition for noisy generatability in the limit in terms of uniform noisy generatability.

### v2

@@ 3.3 Non-uniform Noise-dependent Generatability
Similar to the characterization of non-uniform generatability in the noise-free setting, we can use uniform noise-dependent generatability to provide sufficiency and necessary conditions for non-uniform noise-dependent generatability.
Lemma 3.6 (Sufficiency for Non-uniform Noise-dependent Generatability).
@@
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\cdots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{i}$ and $\operatorname{NC}_{i}(\mathcal{H}_{i})<\infty$ for all $i\in\mathbb{N}$, then $\mathcal{H}$ is non-uniformly noise-dependent generatable.
To prove Lemma 3.6, we construct a non-uniform noise-dependent generator $\mathcal{G}$, which at each round $t\in\mathbb{N}$, computes an index $i_{t}\in\mathbb{N}$ based on the number of distinct examples seen in the stream so far. $\mathcal{G}$ then computes the noisy closure of $\mathcal{H}_{i_{t}}$ at noise-level $i_{t}$ and then plays from this set. We defer details to Appendix F.
As a corollary of Lemma 3.6 and Corollary 3.4, we can show that all countable classes are non-uniformly noise-dependent generatable, and hence noisily generatable in the limit. In some sense, this result shows that amongst countable classes, noisy generation comes for free!
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If $\mathcal{H}$ is countable, then $\mathcal{H}$ is non-uniformly noise-dependent generatable, and therefore also noisily generatable in the limit.
@@
Since non-uniform noise-dependent generatability implies noisy generatability in the limit, it suffices to show that every countable $\mathcal{H}$ is non-uniformly noise-dependent generatable. Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any countable class satisfying the UUS property. Let $h_{1},h_{2},\dots$ be some fixed enumeration of $\mathcal{H}$ and define $\mathcal{H}_{n}=\{h_{i}:i\leq n\}$ for all $n\in\mathbb{N}$. Then, observe that $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\dots$ and that $\mathcal{H}=\bigcup_{n=1}^{\infty}\mathcal{H}_{n}$. Next, observe that $|\mathcal{H}_{n}|=n<\infty$, which, using Corollary 3.4, implies that $\operatorname{NC}_{n}(\mathcal{H}_{n})<\infty$. Finally, Lemma 3.6 gives that $\mathcal{H}$ is non-uniformly noise-dependent generatable. ∎
Corollary 3.7 with Lemma 3.5 also establishes that uniform noise-dependent generatability is strictly harder than non-uniform noise-dependent generatability. We now move to our necessity condition for non-uniform noise-dependent generatability.
Lemma 3.8 (Necessity for Non-uniform Noise-dependent Generatability).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If $\mathcal{H}$ is non-uniformly noise-dependent generatable, then for every $n\in\mathbb{N}$, there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\cdots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{i}$ and $\operatorname{NC}_{n}(\mathcal{H}_{i})<\infty$ for all $i\in\mathbb{N}.$
@@
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any non-uniformly noise-dependent generatable class satisfying the $\operatorname{UUS}$ property. Let $\mathcal{G}$ be a non-uniform noise-dependent generator for $\mathcal{H}$. Fix $n\in\mathbb{N}$. For every $h\in\mathcal{H}$, let $d_{h,n}\in\mathbb{N}$ be the smallest natural number such that for any sequence $x_{1},x_{2},\dots$ with $\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq n,$ if there exists a $t\in\mathbb{N}$ such that |$|\{x_{1},\dots,x_{t}\}|=d_{h,n}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t$. Let $\mathcal{H}_{i}=\{h\in\mathcal{H}:d_{h,n}\leq i\}$ for all $i\in\mathbb{N}$. Note that $\mathcal{H}=\bigcup_{i\in\mathbb{N}}\mathcal{H}_{i}$ because $d_{h,n}<\infty$ for all $h\in\mathcal{H}.$ Moreover, we have that $\mathcal{H}_{i}\subseteq\mathcal{H}_{i+1}$ for all $i\in\mathbb{N}.$ Finally, for every $i\in\mathbb{N}$, observe that $\mathcal{G}$ is a uniform generator for $\mathcal{H}_{i}$ at noise-level $n\in\mathbb{N}$, implying that $\operatorname{NC}_{n}(\mathcal{H}_{i})<i<\infty.$
Note the sufficiency and necessary conditions in Lemmas 3.6 and 3.8 respectively are not matching. Indeed, the condition in Lemma 3.6 is stronger than that in Lemma 3.8. We leave as an open question a complete characterization of non-uniform noise-dependent generatability.
@@ 3.4 Noisy Generatability in the Limit
Corollary 3.7 showed that all countable classes are noisily generatable in the limit. Here, we provide alternate sufficiency conditions for noisy generatability in the limit. Our first result shows that (noiseless) non-uniform generatability is sufficient for noisy generatability in the limit. Since all countable classes are (noiseless) non-uniform generatable (Li et al., 2024), this result also shows that all countable classes are noisily generatable in the limit.
@@
  for $t=1,2,\dots$ do
     Adversary reveals example $x_{t}$
@@
     Let $d_{t}=|\{x_{1},\dots,x_{t}\}|$ and $r_{t}\leq t$ be the largest time point such that |$|\{x_{r_{t}},\dots,x_{t}\}|=\lfloor{\frac{d_{t}}{2}}\rfloor$
     Initialize $\hat{z}^{t}_{1}=\mathcal{Q}(x_{r_{t}:t})$
     for $i=1,\dots,r_{t}-1$ do
        if $\hat{z}^{t}_{i}\notin\{x_{1},\dots,x_{t}\}$ then
           $\mathcal{G}$ outputs $\hat{z}^{t}_{i}$ and moves to next round.
        else
           Update $\hat{z}^{t}_{i+1}=\mathcal{Q}(x_{r_{t}},\dots,x_{t},\hat{z}^{t}_{1},\dots,\hat{z}^{t}_{i})$
        end if
     end for
     $\mathcal{G}$ outputs $\hat{z}^{t}_{r_{t}}$ and moves to next round.
  end for
@@
We will show that $\mathcal{G}$ noisily generates from $\mathcal{H}$ in the limit. Let $h^{\star}\in\mathcal{H}$ and $x_{1},x_{2},\dots$ be the chosen hypothesis and noisy enumeration of $\operatorname{supp}(h^{\star})$ picked by the adversary, respectively. Let $d_{\mathcal{Q}}(h^{\star})\in\mathbb{N}$ be the number of noiseless distinct examples that $\mathcal{Q}$ needs for $h^{\star}$ to perfectly generate from $\operatorname{supp}(h^{\star})$. Let $t^{\star}\in\mathbb{N}$ be such that for all $t\geq t^{\star}$, we have that $x_{t}\in\operatorname{supp}(h^{\star}).$ Such a $t^{\star}$ must exist because there are at most a finite number of noisy examples in the stream. Also, because $x_{1},x_{2},\dots$ is a noisy enumeration, at some point $s^{\star}\in\mathbb{N}$, $r_{s^{\star}}\geq t^{\star}$ and $\lfloor\frac{d_{s^{\star}}}{2}\rfloor\geq d_{\mathcal{Q}}(h^{\star})$. Fix some $s\geq s^{\star}.$ We will prove that $\mathcal{G}$ generates perfectly on round $s$. First, we claim that for all $i\in[r_{s}]$, we have that $\hat{z}^{s}_{i}\in\operatorname{supp}(h^{\star})\setminus\{x_{r_{s}},\dots,x_{s},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{i-1}\}$. Our proof will be by induction. For the base case, consider $i=1$. We need to show that $\hat{z}^{s}_{1}\in\operatorname{supp}(h^{\star})\setminus\{x_{r_{s}},\dots,x_{s}\}$. However, this just follows from the fact that $\hat{z}^{s}_{1}=\mathcal{Q}(x_{r_{s}:s})$, |$|\{x_{r_{s}},\dots,x_{s}\}|\geq d_{\mathcal{Q}}(h^{\star})$, and $\{x_{r_{s}},\dots,x_{s}\}\subseteq\operatorname{supp}(h^{\star}).$ Next, for the induction step, suppose that for all $i\leq m$, we have that $\hat{z}^{s}_{i}\in\operatorname{supp}(h^{\star})\setminus\{x_{r_{s}},\dots,x_{s},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{i-1}\}$. Then, by definition, we have that $\hat{z}_{m+1}^{s}=\mathcal{Q}(x_{r_{s}},\dots,x_{s},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{m})$. The proof of the claim is complete after noting that $\{x_{r_{s}},\dots,x_{s},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{m}\}\subseteq\operatorname{supp}(h^{\star})$ and |$|\{x_{r_{s}},\dots,x_{s},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{m}\}|\geq d_{\mathcal{Q}}(h^{\star}).$
@@
Now we will complete the overall proof by showing that the output of $\mathcal{G}$ on round $s$ lies in $\operatorname{supp}(h^{\star})\setminus\{x_{1},\dots,x_{s}\}$. Suppose that $\mathcal{G}$ outputs $\hat{z}_{j}^{s}$ for some $j\in[r_{s}]$. If $j<r_{s}$, then it must be the case that Line 7 fired and so by construction $\hat{z}_{j}^{s}\in\operatorname{supp}(h^{\star})\setminus\{x_{1},\dots,x_{s}\}$. Thus, suppose that $j=r_{s}.$ If $j=r_{s}$, then it must be the case that $\{\hat{z}_{1}^{s},\dots,\hat{z}_{r_{s}-1}^{s}\}=\{x_{1},\dots,x_{r_{s}-1}\}.$ Accordingly, we have that $\{x_{r_{s}},\dots,x_{s},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{r_{s}-1}\}=\{x_{1},\dots,x_{s}\}$, which means that $\hat{z}^{s}_{r_{s}}\in\operatorname{supp}(h^{\star})\setminus\{x_{r_{s}},\dots,x_{s},\hat{z}^{s}_{1},\dots,\hat{z}^{s}_{r_{s}-1}\}$ and therefore $\hat{z}^{s}_{r_{s}}\in\operatorname{supp}(h^{\star})\setminus\{x_{1},\dots,x_{s}\}.$ Thus, in either case, $\mathcal{G}$ perfectly generates on round $s$. Since $s\geq s^{\star}$ is arbitrary, our proof is complete. ∎
@@
In similar spirit to Theorem 3.10 from Li et al. (2024), we provide another sufficiency condition for noisy generatability in the limit in terms of uniform noise-independent generatability.